Summary:

Another productive night. We struggled early again, due the recovery from a power outage. Otherwise, telescope and camera did well. We did spend fair amount of time in the second half to chase down what turned out to be a Y band focus offset issue which resulted in out-of-focus images. The other persistent problem was blue guidinng which kept on giving trailed images.

Two milestones for tonight: after fixing Y focus, we got 0.5" images in Y. And we were able to track the Bootes field all the way across transit at elevation 88.6 deg.

As for data collection, we did five OBs for UMaII dwarf galaxies under ~1.0" seeing, a long dither sequence for NGC 3348 under 0.9" seeing, and four OBs for Bootes under 0.6-1.0" seeing.

Details

Observers: Xiaohui Fan, Linhua Jiang, Dave Sand, St\xE9phane Herbert-Fort, Eric Nielson

Supporting Astronomer: Olga Kuhn

Operator David Gonzalez Huerta

All Times UT

0000 some low clouds, not going to be photometric.

0006 lost power; we are on generator power. 25Bias_Bino.ob

0110 opening. SkyFlat_UsY_5.ob Us: 20000; Y: 20000; SkyFlat_UsY_rot180_5.ob

0136 SkyFlat_BR_5.ob B: 20000, R: 30000

0147 SkyFlat_Vr_5.ob V: 15000, r: 18000 (only two are done, we need to switch back to commerical power).

0450 after a lot of struggle with both telescope and cameras, we are finally on the first science target

0455 UMaII_pos3_Br.ob B: 1.2", r: 1.0". B image looks a bit trailed in the last few images.

0540 UMaII_pos4_Br.ob seeing 1.1"

0621 UMaII_pos5_Br.ob first two blue images trailed, likely due to guider oscillation. r: 1.0"

0700 made changes to the guiding tolerance from 0.06" to 0.20" to minimize blue movements, as the enlongation could be caused by jumps in the blue. turn off and restart camera.

0710 UMaII_pos6_Br.ob 1.0" on both sides. blue still trails somewhat. John Hill tries to stop blue guiding and let red continue guiding. didn't work. But at least some of the images are good.

0747 focusing. UMII_pos7_Br.ob B: 1.0 (still slightly trailed), r: 0.8"

0833 going to NGC_3368_10step, use blue only. 0.8" in V, still slightly trailed...

0820 going to Bootes field. focusing...

0829 starting b34r, 5x240sec, Y: 0.9", U: 1.2"

0900 starting b41r image looks donut. try fastextra again. seems to have a lot of coma.

0915 starting b41r again, image still looks donuts.

0939 restarting camera and clearing active optics, try again to see if image will improve.

1140 we finally find that the focus offset of Y is off by a lot. The best focus we determined was 0.02, while it was set to be 0.09. At 0.02, we have 0.5" seeing! After correcting focus, we have 240sec image of 0.6". seems only four images here. We have been ignoring U for now. There seems to be a lot of guiding issues, but as we are close to twilight, have to press on...

1205 starting b43r at elevation of 83 deg... 0.5-0.7" images... still going at 87 deg. 4th image looks bad, we are probably out of collimation now.

1238 transit at 88.6 elevation.

1241 b39r first image trailed. 0.7" images. into twilight but Y band is still dark. take another position 1 in the end. sky level still ok.

1325 SkyFlat_Vr_5.ob

1335 SkyFlat_Br_5.ob

-- XiaohuiFan - 08 Feb 2008
Topic revision: r8 - 11 Feb 2008, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback